Operations Report – April 19th

Crew 245 Operations Report 19-04-2021

SOL: 9

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7.5

From what time last night: 7:00 pm, SOC 53%

To what time this morning: 2:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 71% at 7:03 pm (Before generator is run at night)

Notes on power system: For the last four sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (around 7:00 pm compared to well after sunset) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 15 gallons

Water Meter: 150567.2 units

Water (static tank): 265 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 203.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.25 gallon (estimated)

# Hours the ATVs were used today: 4 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for the duration of mission. All ATVs have been re-fueled, and there are 2 gallons of ethanol-free gasoline on campus.

HabCar used and why, where?: Yes, to jump-start the CrewCar battery.

CrewCar used and why, where?: Not used, but the battery was dead, so it was jump-started.

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiments. The Crew Botanist briefly checked plant growth and health in GreenHab today. Used briefly for crew documentary filming.

Summary of ScienceDome operations: Used briefly for crew documentary filming.

Summary of RAM operations: Used briefly for crew documentary filming.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Sol Summary #9 – April 19th

Crew 245 Sol Summary Report 19-04-2021

Sol: 9

Summary Title: The Day of “Rest”

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 09:00-11:00: Pancakes and hanging out while slowly briefing for what’s needed today.

· 11:30-13:30: Crew split to work on different items around the Hab.

· 13:30-14:30: Lunch together!

· 14:30-15:30: Working on separate projects around the Hab.

· 15:30-17:30: Recording all footage/photos needed for STEM outreach and for our mockumentary!

· 17:30-19:00: Dinner! We recorded dinner as we cooked it this time.

· 19:00 on: Free time? Weird…

Look Ahead Plan: Tomorrow we have an EVA to the north region for the first time for crew #2!

Anomalies in work: None.

Weather: Windy! We stayed in for a good reason.

Crew Physical Status: No issues.

EVA: None

Reports to be filed: Operations Report, EVA Request

Support Requested: n/a